### v1

# ArMeme: Propagandistic Content in Arabic Memes

Abstract
With the rise of digital communication memes have become a significant medium for cultural and political expression that is often used to mislead audience. Identification of such misleading and persuasive multimodal content become more important among various stakeholders, including social media platforms, policymakers, and the broader society as they often cause harm to the individuals, organizations and/or society. While there has been effort to develop AI based automatic system for resource rich languages (e.g., English), it is relatively little to none for medium to low resource languages. In this study, we focused on developing an Arabic memes dataset with manual annotations of propagandistic content. [^1] Propaganda is a form of communication designed to influence people’s opinions or actions toward a specific goal, employing well-defined rhetorical and psychological techniquesfor Propaganda Analysis (1938). We annotated $\sim 6K$ Arabic memes collected from various social media platforms, which is a first resource for Arabic multimodal research. We provide a comprehensive analysis aiming to develop computational tools for their detection. We will make them publicly available for the community.
ArMeme: Propagandistic Content in Arabic Memes
Firoj Alam1, Abul Hasnat2,3, Fatema Ahmed1, Md Arid Hasan4, Maram Hasanain1 1Qatar Computing Research Institute, HBKU, Qatar 2Blackbird.AI, USA, 3APAVI.AI, France 4University of New Brunswick, Fredericton, NB, Canada {fialam,mhasanain,fakter}@hbku.edu.qa, mhasnat@gmail.com, arid.hasan@unb.ca
1 Introduction
Social media platforms have enabled people to post and share content online. A significant portion of this content provides valuable resources for initiatives such as citizen journalism, raising public awareness, and supporting political campaigns. However, a considerable amount is posted and shared to mislead social media users and to achieve social, economic, or political agendas. In addition, the freedom to post and share content online has facilitated negative uses, leading to an increase in online hostility, as evidenced by the spread of disinformation, hate speech, propaganda, and cyberbullying Brooke (2019); Joksimovic et al. (2019); Schmidt and Wiegand (2017); Davidson et al. (2017); Da San Martino et al. (2019a); Van Hee et al. (2015). A lack of media literacy [^2] Media literacy encompasses the ability to access, analyze, evaluate, and create media in various forms. is also a major factor contributing to the spread of misleading information on social media Zannu et al. (2024). This can lead to the uncritical acceptance and sharing of false or misleading content, which can quickly disseminate through social networks. In their study, Zannu et al. (2024) highlight the crucial role of media literacy in mitigating the spread of fake news among users of platforms such as Instagram and Twitter.
Figure 1: Examples of Arabic memes representing different categories.
Online content typically consists of different modalities, including text, images, and videos. Disinformation, misinformation, propaganda, and other harmful content are shared across all these modalities. Recently, the use of Internet memes have become very popular on these platforms. A meme is defined as “a collection of digital items that share common characteristics in content, form, or stance, which are created through association and widely circulated, imitated, or transformed over the Internet by numerous users” Shifman (2013). Memes typically consist of one or more images accompanied by textual content Shifman (2013); Suryawanshi et al. (2020). While memes are primarily intended for humor, they can also convey persuasive narratives or content that may mislead audiences. To automatically identify such content, research efforts have focused on addressing offensive material Gandhi et al. (2020), identifying hate speech across different modalities Gomez et al. (2020); Wu and Bhandary (2020), and detecting propaganda techniques in memes Dimitrov et al. (2021a).
Among the various types of misleading and harmful content, the spread of propagandistic content can significantly distort public perception and hinder informed decision-making. To address this challenge, research efforts have been specifically directed towards defining techniques and tackling the issue in different types of content, including news articles Da San Martino et al. (2019), tweets Alam et al. (2022b), memes Dimitrov et al. (2021a), and textual content in multiple languages Piskorski et al. (2023a). Most of these efforts have focused on English, with relatively little attention given to Arabic. Prior research on Arabic textual content includes studies presented at WANLP-2022 and ArabicNLP-2023 Alam et al. (2022b); Hasanain et al. (2023). However, for multimodal content, specifically memes, there are no available datasets or resources. To address this gap, we have collected and annotated a dataset consisting of approximately 6,000 memes, categorizing them into four categories (as shown in Figure 1) to identify propagandistic content. Below we briefly summarize the contribution of our work.
  • The first Arabic meme dataset with manual annotations defining four categories.
  • A detailed description of the data collection procedure, which can assist the community in future data collection efforts.
  • An annotation guideline that will serve as a foundation for future research.
  • Detailed experimental results, including:
    • Text modality: training classical models and fine-tuning monolingual vs. multilingual transformer models.
    • Image modality: fine-tuning CNN models with different architectures.
    • Multimodality: training an early fusion-based model.
    • Evaluating different LLMs in a zero-shot setup for all modalities.
  • Releasing the dataset to the community. [^3] Please contact the authors for early access to the dataset. The dataset and annotation guideline will be beneficial for research to develop automatic systems and enhance media literacy.
2 Related Work
The widespread use of social media has become one of the main ways of sharing information and is also responsible for creating and spreading misinformation and propaganda among users. Propagandistic techniques often utilize various types of content, such as fake news and doctored images, across multiple media platforms, frequently employing tools like bots. This information is distributed in diverse forms, including textual, visual, and multi-modal. To mitigate the impact of propaganda in online media, researchers have been developing resources and tools to identify and debunk such content.
2.1 Persuasion Techniques Detection
Early research on propaganda identification relies on the entire document to identify whether the content is propaganda, while recent studies focus on social media content Dimitrov et al. (2021b), news articles Da San Martino et al. (2019b), political speech Partington and Taylor (2017), arguments Habernal et al. (2017, 2018), and multimodal content Dimitrov et al. (2021a). Barrón-Cedeno et al. (2019) developed a binary classification (propaganda and non-propaganda) corpus to explore writing style and readability levels. An alternative approach followed by Habernal et al. (2017, 2018) to identify persuasion techniques within the texts constructing a corpus on arguments. Moreover, the study of Da San Martino et al. (2019b) developed a span-level propaganda detection corpus from news articles and annotated in eighteen propaganda techniques.
Piskorski et al. (2023b) developed a dataset from online news articles into twenty-two persuasion techniques containing nine languages to address the multilingual research gap. Following the previous work, Piskorski et al. (2023a) and SemEval-2024 task 4 focus on resource development to facilitate the detection of multilingual persuasion techniques. Focusing on multimodal persuasion techniques for memes, Dimitrov et al. (2021a) created a corpus containing 950 memes and investigated pretrained models for both unimodal and multimodal memes. The study of Chen et al. (2024) proposed a multimodal visual-textual object graph attention network to detect persuasion techniques from multimodal content using the dataset described in Piskorski et al. (2023b). In a recent shared task, Dimitrov et al. (2024) introduced a multilingual and multimodal propaganda detection task, which attracted many participants. The participants’ systems included various models based on transformers, CNNs, and LLMs.
2.2 Multimodal Content
The study of multimodal content has gained popularity among researchers for propaganda detection due to the effectiveness of multimodal content in spreading propaganda information and creating positive impacts among the targeted audience. Sharma et al. (2022) presented propaganda can be used to cause several types of harm including hate, violence, exploitation, etc. while spreading mis- and dis-information is also one of the main reasons Alam et al. (2022a). The study of Volkova et al. (2019) presented an in-depth analysis of multimodal content for predicting misleading information from news. Additionally, the deception and disinformation analysis on social media platforms using multimodal content in multilingual settings has been studied by Glenski et al. (2019). Moreover, hateful memes Kiela et al. (2020), propaganda in visual content Seo (2014), emotions and propaganda Abd Kadir et al. (2016) also studied by the researchers in the past few years.
Recent studies focusing on fine-tuning visual transformer models such as ViLBERT Lu et al. (2019), Multimodal Bitransformers Kiela et al. (2019), and VisualBERT Li et al. (2019). Cao et al. (2022) study focuses on multimodal hateful meme identification using prompting strategies by adopting Prakash et al. (2023). Hee et al. (2024) studied hate speech content moderation and discussed recent advancements leveraging large models.
Compared to previous studies, our work differs in that we provide the first resource for Arabic. Additionally, our annotation guidelines and data collection procedures for memes may be useful for other languages.
3 Dataset
Figure 2: Data curation pipeline.
Our data collection process involve several steps as highlighted in the Figure 2. We manually selected public groups from Facebook, Instagram, and Pinterest. In addition, we have also collected memes from Twitter using a set of keywords as listed in the Figure  3 (in Appendix). Our data curation consists of a series of steps as discussed below.
Manual selection of groups, links and keywords:
Focusing on the mentioned sources we have manually selected public groups, which contains post on public figures, celebrity, and mentions about politics.
Crawling:
Given that Facebook, Instagram and Pinterest do provide API or do not allow automatic crawling images, therefore, we developed a semi-automatic approach to crawl images from these platforms. The steps include manually loading images and then crawl the images that are loaded on the browser. For the Twitter (X-platform), we used the keywords to crawl tweets, which consists of media/image.
Filtering duplicate images:
Given that user might have posted same meme or a slight modification of it in multiple platforms, which is very common for social media, therefore, we applied an exact and near-duplicate image detection method to remove them. This method consists of extracting features using a pre-trained deep learning model and compute similarity. Given a dataset $\mathcal{D}=\{x_{1},x_{2},\ldots,x_{N}\}$ consisting of $N$ data points, we extracted features using a pre-trained deep learning model and used nearest neighbor based approach Cunningham and Delany (2007). The model is trained by fine-tuning ResNet18 He et al. (2016) using the social media dataset discussed in Alam et al. (2020). Let $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ be a pre-trained deep learning model that maps an input data point $x_{i}\in\mathbb{R}^{d}$ to a feature vector $f(x_{i})\in\mathbb{R}^{m}$. For each data point $x_{i}\in\mathcal{D}$, the feature vector is extracted as: $\mathbf{z}_{i}=f(x_{i}),\quad\text{for }i=1,2,\ldots,N$ where $\mathbf{z}_{i}\in\mathbb{R}^{m}$ is the feature vector of the data point $x_{i}$. To compute the nearest neighbors between a data point $x_{i}$ and the entire dataset $\mathcal{D}$, we use the euclidean distance. We then use a threshold of $3.6$ to define the near-duplicate images as those with a euclidean distance less than or equal to this threshold value.
OCR Text:
We used EasyOCR [^4] https://github.com/JaidedAI/EasyOCR to extract text from memes. Memes with no extracted OCR text were filtered out.
Classifier-Based Filtering:
We employed an in-house meme vs. non-meme classifier to filter out images that were not classified as memes. The classifier was developed using a dataset of 3,935 images, consisting of 2,000 memes and 1,935 non-memes. Following the approach of Hasnat et al. (2019), we developed a lightweight meme classifier to perform binary classification based on the extracted image features. The classifier achieved the best performance of 94.79% test set accuracy in classifying memes using a 256-dimensional normalized histogram extracted from gray-scale images as features, with a Multilayer Perceptron (MLP) as the classifier.
Data Sampling:
Due to budget constraints for manual annotation, we randomly sampled $\sim$6K images.
Manual Annotation:
For the manual annotation, we first prepared an annotation guideline to assist the annotators. The annotation tasks consist of two phases: (1) Phase 1 (meme categorization): labeling memes as (i) not-meme, (ii) other, (iii) not propaganda, or (iv) propaganda; (2) Phase 2 (text editing): editing the text to fix OCR errors. To facilitate the annotation tasks, we developed an annotation platform as presented in Appendix D. The details of the annotation guidelines are reported in Appendix C. Note that we developed the annotation guidelines in English, which were then translated into the Arabic language as detailed in Section (Appendix Section E). In phase 1 (meme categorization), each meme was annotated by three annotators and final label is decided based on majority agreement.
In Figure 1, we provide examples of memes representing different categories. Figure 1(a) depicts a couple in what appears to be a therapy session. The therapist asks, “Do you feel your wife is controlling you?” The wife responds, “No, I don’t feel so.” It is evident that the question was directed towards the husband, yet the wife answers instead of him. The irony lies in her controlling the conversation when her control is the subject of discussion. This meme attempts to humorously portray the stereotypical notion that wives are controlling in marriages. Figure 1(b) employs a play on words to create humor but does not contain any propagandistic techniques. Figure 1(c) features a meme that uses an image of a scene with dialogue and added text to create humor. However, it was categorized as “other” because the dialogues were in English, rather than “not propagandistic” or “propagandistic.” Figure 1(d) shows a picture of book covers, which might have been part of an advertisement.
Annotation Agreement
For the Phase 1 annotation, we computed annotation agreement using various evaluation measures, including Fleiss’ kappa, Krippendorff’s alpha, average observed agreement, and majority agreement. The resulting scores were 0.529, 0.528, 0.755, and 0.873, respectively. Based on the value of Krippendorff’s alpha, we can conclude that our annotation agreement score indicates moderate agreement. [^5] Note that Kappa values of 0.21–0.40, 0.41–0.60, 0.61–0.80, and 0.81–1.0 correspond to fair, moderate, substantial, and perfect agreement, respectively Landis and Koch (1977). In the final label selection, we excluded the $\sim$200 memes on which the annotators disagreed. In the second phase, we mainly edited text to fix the OCR errors, which has been done by a single annotator. To ensure the quality of the editing phase a random samples were checked by an expert annotator and periodically provided feedback. Note that the post-editing has been done for only propagandistic and not-propagandistic memes. It is to reduce the cost of the annotation and to further annotate them with span-level propaganda technique annotation.
Statistics:
Table 1 shows the number of memes for each category. For the rest of the experiments, the data was split into train, dev, and test as shown in the table. The dataset comprises a total of 5,725 annotated samples, with “Not propaganda” covers over half of the dataset ($\sim$66%), followed by “Propaganda.” The “Not-meme“ and “Other“ classes are significantly smaller in comparison. The distribution indicates a significant class imbalance, particularly between “Not propaganda” and the other classes, which could affect model training and performance.
In Table 2, we report the distribution of the dataset across different sources. The annotated number of memes reflects the memes we collected from various sources, as detailed in Table 5. We have the highest number of memes collected and annotated from Instagram. A very small number from Twitter is due to different image filtering steps. As shown in Table 2 the prevalence of propagandistic memes is relatively higher on Facebook than that of non-propagandistic memes.
Class label Train Dev Test Total
Not propaganda 2,634 384 746 3,764
Propaganda 972 141 275 1,388
Not-meme 199 30 57 286
Other 202 29 56 287
Total 4,007 584 1,134 5,725
Table 1: Data split statistics.
Source Not prop. Prop. Not-meme Other Total
Facebook 464 332 58 144 998
Instagram 2,052 637 46 60 2,795
Pinterest 1,245 414 147 78 1,884
Twitter 3 5 38 2 48
Total 3,764 1,388 289 284 5,725
Table 2: Number of annotated memes across different sources. Prop. - Propaganda.
4 Experiments
4.1 Training and Evaluation Setup
For all experiments, except for those involving LLMs as detailed below, we trained the models using the training set, fine-tuned the parameters with the development set, and assessed their performance on the test set. We use the model with the best weighted-F1 on the development set to evaluate its performance on the test set. For the LLMs, we accessed them through APIs.
Evaluation Measures
For the performance measure for all different experimental settings, we compute accuracy, and weighted precision, recall and F1 score. In addition, we also computed macro-F1.
4.2 Models
We conducted our experiments using classical models (e.g., SVM) as well as both small (e.g., ConvNeXt-T) and large language models. It is important to note that our definitions of ‘small’ and ‘large’ models are based on the criteria discussed in Zhao et al. (2023). [^6] The term ‘LLMs’ specifically refers to models that encompass tens or hundreds of billions of parameters.
4.2.1 Baseline:
We adopted widely-used standard baseline methods, including the majority and random baselines.
4.2.2 Small Language Models (SLMs)
We implemented classical models across all modalities, consisting of (i) feature extraction followed by model training, and (ii) fine-tuning pre-trained models (PLMs). For fine-tuning PLMs, we used a task-specific classification head over the training subset.
Text-Based Models:
For the text-based unimodal model, we transformed text into $n$-gram (n=1) format using a tf-idf representation, considering the top 5,000 tokens, and trained an SVM model with a parameter value of $C=1$. Additionally, we fine-tuned several pre-trained transformer models (PLMs). These included the monolingual transformer model AraBERT (Antoun et al., 2020), Qarib Abdelali et al. (2021) and multilingual transformers such as multilingual BERT (mBERT) Devlin et al. (2019), and XLM-RoBERTa (XLM-r) (Conneau et al., 2019). We used the Transformer toolkit Wolf et al. (2019) for the experiment. Following the guidelines outlined in (Devlin et al., 2019), we fine-tuned each model using the default settings over three epochs. Due to instability, we performed ten reruns for each experiment using different random seeds, and we picked the model that performed best on the development set. We provided the details of the parameters settings in Appendix B.
Image-Based Models:
For the image-based unimodal model with feature-extraction approach, we extracted features using ConvNeXt-T Liu et al. (2022), [^7] The configuration of ConvNeXt-T includes $C=(96,192,384,768)$ and $B=(3,3,9,3)$, where $C$ and $B$ represent the number of channels and blocks, respectively. and trained an SVM model. For fine-tuning image-based PLMs, we used ResNet18, ResNet50 He et al. (2016), VGG16 Simonyan and Zisserman (2014), MobileNet Howard et al. (2017), and EfficientNet Tan and Le (2019). We chose these diverse architectures to understand their relative performance. The models were trained using the Adam optimizer Kingma and Ba (2015) with an initial learning rate of $10^{-3}$, which was decreased by a factor of 10 when accuracy on the development set stopped improving for 10 epochs. The training lasted for 150 epochs.
Multimodal Models:
We developed a multimodal model by concatenating text features (extracted using AraBERT) and image features (extracted using ConvNeXt-T), which were then fed into an SVM.
4.2.3 LLMs for Text
For the LLMs, we investigate their performance with zero-shot learning settings without any specific training. It involves prompting and post-processing of output to extract the expected content. Therefore, for each task, we experimented with a number of prompts. We used GPT-4 OpenAI (2023). We set the temperatures to zero for all these models to ensure deterministic predictions. We used LLMeBench framework Dalvi et al. (2024) for the experiments, which provides seamless access to the API end-points and followed prompting approach reported in Abdelali et al. (2024).
4.2.4 Multimodal LLMs
For the multimodal models Xu et al. (2023), we experimented with several well-known and top-performing commercial models. These included OpenAI’s GPT models (GPT-4 Turbo and GPT-4o) OpenAI (2023), as well as Google’s Gemini Pro models (versions 1.0 and 1.5) Team et al. (2023).
Using these models, we tested (i) the meme/image only, (ii) text only (text extracted using OCR from the image), and (iii) multimodal (meme and OCR text) in a zero-shot learning setting. This means we did not provide any training examples within the prompts to the models.
We designed a prompt based on trial and error using the visual interfaces of OpenAI’s GPT-4 user interface. The prompt instructs the models to perform a deeper analysis of the image and any text that they can read within the image before answering whether the meme can be classified as spreading propaganda. Additionally, it requests the models to provide the output in a valid JSON format. For the experiments, we used the default parameters for each multimodal model.
4.3 Prompting Strategy
LLMs produce varied responses depending on the prompt design, which is a complex and iterative process that presents challenges due to the unknown representation of information within different LLMs. The instructions expressed in our prompts include English language with the input text content in Arabic.
As mentioned earlier we employed zero-shot prompting, providing natural language instructions that describe the task and specify the expected output. This approach enables the LLMs to construct a context that refines the inference space, yielding a more accurate output. In Listing 1, we provide an example of a zero-shot prompt, emphasizing the instructions and placeholders for both input and label. Along with the instruction we provide the labels to guide the LLMs and provide information on how the LLMs should present their output, aiming to eliminate the need for post-processing.
Instructions:
prompt = (
"You are an expert social media image analyzer specializing in identifying propaganda in Arabic contexts. "
"I will provide you with Arabic memes and the text extracted from these images. Your task is to briefly analyze them. "
"To accurately perform this task, you will: (a) Explicitly focus on the image content to understand the context and provide a meaningful description and "
"(b) pay close attention to the extracted text to enrich your description and support your analysis. "
"Finally, provide response in valid JSON format with two fields with a format: {\"description\": \"text\", \"classification\": \"propaganda\"}. Output only json. "
"The \"description\" should be very short in maximum 100 words and \"classification\" label should be \"propaganda\" or \"not-propaganda\" or \"not-meme\" or \"other\". "
"Note, other is a category, which is used to label the image that does not fall in any of the previous category."
payload = {
"model": "gpt-4o",
"messages": [
"role": "user",
"content": [
"type": "text",
"text": prompt
"type": "image_url",
"image_url": {
"url": f"data:image/jpeg;base64,{base64_image}"
"type": "text",
"text": Text extracted: + t_text
Listing 1: Zero-shot prompt example for GPT-4.
5 Results and Discussion
In Table 3, we report the detailed classification results for different modalities and models. All models outperform the majority and random baselines. Among the text-based models, the fine-tuned Qarib model outperforms all other models, achieving the best results (0.690 weighted F1) across all modalities and models. AraBERT is the second-best fine-tuned model, with a weighted F1-score of 0.666 among the text-based models. The performance of multilingual transformer models is relatively worse than that of monolingual models.
For the image-based models, the fine-tuned ResNet50 shows the best result (0.673 weighted F1) among all other fine-tuned models and GPT-4o model. The performance of MobileNet (v2) and CNeXt + SVM rank as the second and third best among the fine-tuned models. The results of VGG16 and EfficientNet (b7) are almost similar.
For the multimodal models, the model trained with ConvNeXt + AraBERT + SVM shows the highest performance (0.659 weighted F1) among the multimodal LLMs. The performance of Gemini is significantly worse than that of the GPT-4 variants. GPT-4o demonstrates higher performance compared to GPT-4 Vision.
In our experiments all multimodal model are tested using zero-shot setting, therefore, such lower performance compared to the fine-tuned models are expected.
Model Acc W-P W-R W-F1 M-F1
Baseline
Majority 0.6578 0.4328 0.6578 0.5221 0.1984
Random 0.4797 0.5184 0.4797 0.4797 0.2397
Unimodal - Text
Ngram 0.669 0.624 0.669 0.582 0.280
AraBERT 0.688 0.670 0.688 0.666 0.511
Qarib 0.697 0.688 0.697 0.690 0.551
mBERT 0.707 0.688 0.707 0.675 0.487
XLM-r 0.699 0.676 0.699 0.678 0.489
GPT-4v 0.664 0.620 0.664 0.624 0.384
GPT-4o 0.573 0.611 0.573 0.579 0.350
Unimodal - Image
CNeXt + SVM 0.655 0.608 0.655 0.614 0.405
MobileNet (v2) 0.660 0.618 0.660 0.620 0.426
ResNet18 0.656 0.597 0.656 0.593 0.358
ResNet50 0.660 0.638 0.660 0.637 0.434
Vgg16 0.656 0.597 0.656 0.593 0.358
Eff (b7) 0.660 0.597 0.660 0.595 0.352
GPT-4v 0.565 0.551 0.565 0.545 0.223
GPT-4o 0.693 0.627 0.693 0.634 0.305
Multimodal
CNeXt + ArB + SVM 0.683 0.655 0.683 0.659 0.513
Gemini 0.519 0.551 0.519 0.521 0.276
GPT-4v 0.681 0.461 0.330 0.619 0.340
GPT-4o 0.653 0.443 0.354 0.639 0.363
Table 3: Classification with different modalities. CNeXt: ConvNeXt, Eff (b7): Efficientnet (b7), Gemini: Gemini-1.5-flash-preview-0514l, GPT-4v: GPT-4-vision (gpt-4-vision-preview) W-*: weighted average; M-: Macro average. XLM-r: XLM-RoBERTa base.
6 Additional Experiments
We further conducted experiments using the dataset released as part of the ArAIEval shared task 2 Hasanain et al. (2024), focusing on two labels: propaganda and not-propaganda. The dataset statistics are provided in Appendix A.2. The goal was to investigate model performance in a binary classification scenario.
Table 4 presents the competitive results of four multimodal models with image-only input: GPT-4o, GPT-4 Turbo, and Gemini Pro 1.0. Among these models, GPT-4o significantly outperforms the others and demonstrates the highest performance across all evaluated metrics, achieving an accuracy of 85.17%, a precision of 84.80, a recall of 85.17, and a weighted F1-score of 84.87. In comparison, GPT-4 Turbo lags behind GPT-4o in all metrics, with an accuracy of 76.44%, indicating a significant performance drop compared to GPT-4o. Gemini Pro 1.0 shows lower performance than the GPT-4 models, with an accuracy of 72.47%.
Model Acc. W-P W-R W-F1 M-F1
Gemini 0.725 0.685 0.725 0.663 0.345
GPT-4v 0.764 0.748 0.764 0.735 0.645
GPT-4o 0.852 0.848 0.852 0.849 0.810
Table 4: Results on ArAIEval dataset. Gemini: version Pro 1.0.
7 Conclusions and Future Work
In this study, we introduce a manually annotated dataset for detecting propaganda in Arabic memes. We have annotated $\sim$ 6K memes with four different categories, making it the first such resource for Arabic content. To facilitate future annotation efforts for this type of content, we developed annotation guidelines in both English and Arabic and are releasing them to the community. Our work provides an in-depth analysis of the dataset and includes extensive experiments focusing on different modalities and models, including pre-trained language models (PLMs), large language models (LLMs), and multimodal LLMs. Our results indicate that fine-tuned models significantly outperform LLMs.
In future work, we plan to extend the dataset with further annotations that include hateful, offensive, and propagandistic techniques.
8 Limitations
The dataset we have collected originates from various public groups on Facebook, Instagram, Pinterest, and Twitter. The annotated dataset is highly imbalanced, which may affect model performance. Therefore, it is important to develop models with this aspect in mind.
Ethics and Broader Impact
Our dataset solely comprises memes, and we have not collected any user information; therefore, the privacy risk is nonexistent. It is important to note that annotations are subjective, which inevitably introduces biases into our dataset. However, our clear annotation schema and instructions aim to minimize these biases. We urge researchers and users of this dataset to remain critical of its potential limitations when developing models or conducting further research. Models developed using this dataset could be invaluable to fact-checkers, journalists, and social media platforms.
Acknowledgments
The contributions of F. Alam, M. Hasanain, and F. Ahmed were funded by the NPRP grant 14C-0916-210015, which is provided by the Qatar National Research Fund (a member of Qatar Foundation).
Appendix A Additional Dataset Details
A.1 Data Sources
In Table 5, we provide the sources of the dataset, number of groups and number of image we have collected.
Source # of Group # of Images
Facebook 19 5,453
Instagram 22 107,307
Pinterest - 11,369
Twitter - 5,369
Total 129,498
Table 5: Statistics of the initial data collection.
Figure 3: Keywords used to collect tweets.
A.2 ArAIEval Shared Task Dataset
Table 6 report the distribution of the dataset released as a part of the ArAIEval shared task. For our study we benchmarked this dataset using multimodal models.
Class labels Train Dev Test Total
Not propaganda 1,540 224 436 2,200
Propaganda 603 88 171 862
Total 2,143 312 607 3,062
Table 6: Distribution of dataset for ArAIEval shared task 2.
Appendix B Details of the experiments
For the experiments with transformer models, we adhered to the following hyper-parameters during the fine-tuning process. Additionally, we have released all our scripts for the reproducibility.
  • Batch size: 8;
  • Learning rate (Adam): 2e-5;
  • Number of epochs: 10;
  • Max seq length: 256.
Models and Parameters:
  • AraBERT: L=12, H=768, A=12; the total number of parameters is 371M.
  • XLM-RoBERTa (xlm-roberta-base): L=24, H=1027, A=16; the total number of parameters is 355M.
Appendix C Annotation Task
We designed the annotation instructions through careful analysis and discussion, followed by iterative refinements based on observations and input from the annotators based on the pilot annotation. Our annotation schema is structured into two phases.
C.1 Phases of Annotations
To ensure the quality of the annotation and facilitate the work of annotators, we conducted the annotation in two phases: (i) meme categorization and (ii) text editing. The first phase (see Section C.2) focuses primarily on categorization. In the second phase (see Section C.3), our goal is to edit the text only for memes labeled as propagandistic or not propagandistic. The motivation for editing the text for these categories is to further utilize them for other annotation tasks. For example, propagandistic memes can be further annotated with specific propagandistic techniques. In Figure , we illustrate the thought process of the meme annotation phases.
Figure 4: A visual representation of the annotation process. Block with yellow color represents phase 2.
C.2 Meme Categorization
C.2.1 Definition of a Meme:
Memes typically consist of a background image, which could be a photograph, illustration, or screenshot, and a layer of text that adds context, humor, or commentary to the image. The text is usually placed at the top and/or bottom of the image but not always. The combination of the image and the text creates a specific message, joke, or commentary that is meant to be easily understood, relatable, and shareable. Some characteristics of memes as observed during analysis and discussion:
  1. 1.
    Text overlaid on image.
  2. 2.
    The text has humor in it.
  3. 3.
    The image must meet points 1 and 2.
  4. 4.
    Some contents of the image have been edited.
  5. 5.
    Text might be added to different locations of the image.
  6. 6.
    Mostly uses images of entities with facial expressions (human, animals, fictional characters), which are then used to construct meaning alongside the added text.
  7. 7.
    Uses an entity performing a certain action that might be used to construct meaning alongside the added text.
  8. 8.
    Uses an entity that represents an idea or culture, to construct meaning alongside the added text.
  9. 9.
    Mostly uses screenshots from movie scenes and dialogues with added comments, to create memes.
  10. 10.
    Most of the pictures used to make the meme can be re-edited and a new funny comment can be added to it.
Note: In points 6, 7, and 8, the removal of the entity from the images will affect the meaning. In other words, if the entity is removed, then the meaning will not be complete. This is what we mean by constructing meaning.
C.2.2 Defining Propaganda:
Propaganda is any communication that deliberately misrepresents symbols and/or entities, appealing to emotions and prejudices while bypassing rational thought, to influence its audience toward a specific goal. Memes are created to be humorous; therefore, it is natural that they lack rational discussion. Instead, they use content to appeal to emotions and prejudices. For our task, we defined the following four categories and annotated the memes accordingly.
(1) Not-Meme:
For images that do not follow the definition of a meme, examples of images labeled as “not-meme” are shown in Figure 5.
Figure 5: Examples of images labeled as not-meme.
(2) Other:
For images that can be defined as memes but fall under any of the criteria listed below. Examples of images labeled as “not-meme” are shown in Figure 6. The criteria for "Other":
  1. 1.
    Memes that rely on nudity and offensive content, unless the target of the offense is a famous, political, or religious entity.
  2. 2.
    Memes that rely on numbers or figures to construct meaning.
  3. 3.
    Memes that show explicit nudity.
  4. 4.
    Memes that explicitly use offensive words.
  5. 5.
    Memes that are in a different language (not Arabic).
  6. 6.
    Memes that you could not understand due to the dialect it was written in, poor font size, or for any other reason.
Note: Memes might contain words that have an implicitly offensive meaning, or the use of offensive words may be aimed at social, religious, or political groups. In these cases, the meme does not fall under this criterion.
Figure 6: Examples of images labeled as other.
(3) Not Propaganda:
For memes that follow the definition of memes but do not contain any propaganda techniques, examples of images labeled as “not propagandistic” are shown in Figure 7.
Figure 7: Examples of images labeled as not propaganda.
Figure 8: Examples of images labeled as propaganda.
(4) Propaganda:
For memes that follow the definition of memes and contain propaganda techniques, examples of images labeled as “propagandistic” are shown in Figure 8.
C.3 Text Editing
The task is to edit the text to match the text shown in the image. The interface will show you the picture, alongside the text that is viewed in it. The text was extracted automatically, so it might contain errors. It might not reflect all you see in the picture. Some important guidelines to follow for editing the text are listed below:
  1. 1.
    Each part that is a standalone sentence and makes complete meaning should be written as one line.
  2. 2.
    Punctuation marks are considered a part of the text. They need to be edited/added.
  3. 3.
    If the text is in columns, put first all the text of the first column, then all the text of the next column. This task will specifically address memes in Arabic, so the first column should be considered from the right. However, this is not a rule, and memes might change this orientation, so it is up to the annotator to decide the order based on their understanding.
  4. 4.
    Rearrange the text so that there is one sentence per line, if possible.
  5. 5.
    If there are separate blocks of text in different locations of the image, start a new line from each block.
  6. 6.
    Leave a blank between two blocks of text if they were shown in two different locations on the picture.
  7. 7.
    Items that should be excluded from the text:
    • Usernames and social media account names (if visible in the image).
    • Websites, logos, and any text that is not a part of the meme, so that removing that part does not affect the meaning of the meme.
    • Any text that is hidden and is hard to read.
  8. 8.
    In special cases, a logo can be used in the meme to create meaning. In this case, add the text of the logo to the edited text, if needed.
Example 1:
Figure 9 shows an example of a meme, for editing the text that can be viewed it, the following points are important:
Figure 9: An example of a meme for editing text.
  • Each dialog box is one sentence
  • Start a new line for each box (each box is a different block of text)
  • Remove any elements that are not part of the meaning: account name and location
  • Add or modify punctuation to suit what is presented in the text
  • Text after modification (text translated to EN and read from the first speech bubble from right):
    Get him ... Get him... corner him...
    get him so we can give him his rights
    come... aren’t you coming??
    come...take your rights you son of
    a bastard
    Wallah we gonna get you till...
    we give you all your rights
    you chick ....
Example 2:
Figure 10 shows another example, for which the following points are important.
Figure 10: An example of a meme for editing text.
  • Text written in red is difficult to understand and read, so it should not be included in the text.
  • The text written on the hat and the text in black are each a different block of text. Start a line for each of them and leave a space for each new line.
  • This example is for illustrative purposes only, and “memes” in English will not be shown in this task.
  • Text after modification:
    Bernie
    Riding with Biden **2020**
    Haha hey its the Obama guy
Appendix D Annotation Platform
D.1 Meme Categorization Task
In Figure 11, we provide a screenshot of the annotation platform for the meme categorization task. As shown in the figure, the platform displays the meme itself on the right, the extracted text on the left, a link to the annotation guidelines, and labels with buttons at the bottom for selecting a category for the meme. The task of the annotator was to label the meme as one of the below categories, according to the definitions detailed in the guideline (see Section C). To facilitate the work of annotators in the annotation process, we used the keywords ‘meme’ along with the labels ‘other’, ‘propaganda’, and ‘not-propaganda’.
  • Not Meme
  • Meme, Other
  • Meme, Not Propaganda
  • Meme, Propaganda
Figure 11: A screenshot of the annotation platform for the meme categorization task.
Given that the memes we collected were from different social media platforms, they may contain offensive content. Therefore, we added a note that some pictures may contain offensive content, and that we apologize for any inconvenience that such content may cause. We appreciate your contribution to this project which will minimize the spread of such harmful content on the internet.
To further guide the annotation process, we asked the annotators to follow the following steps.
  1. 1.
    Begin by determining whether the image presented is a “meme”. If the image is not a meme, select “Not Meme”, then click “Submit”. The next image will then be loaded.
  2. 2.
    If the image is a “meme”, assess whether it falls under the category of “Other”. If so, select “Other”, then click “Submit”. The next image will then be loaded.
  3. 3.
    If the image does not fall under the category of “Other”, choose one of the remaining two labels based on your interpretation of the meme’s content. After selecting the appropriate label, edit the text as needed.
D.2 Text Editing Task
In this phase, the task was to edit the text based on the guidelines discussed in Section C.3. In Figure 12, we provide a screenshot demonstrating the text extracted from OCR, an editable text box, and the original meme. The task was to edit the text to match it with the original meme.
Figure 12: An screenshot of the annotation platform for the text editing.